OpenStudy (katielong):

Newton's second law gives m (dv/dt) = mg - cv. Show that the solution to the differential equation is v=(mg/c)(1-e^(ct/m)). Thanks in advance!

OpenStudy (katielong):

Original equation: \[m \frac{ dv }{ dt }=mg-cv\] divided by m and rearranged: \[\frac{ dv }{ dt }+\frac{ cv }{ m }=g\] Integration factor:\[IF=e ^{\int\limits_{}^{}\frac{ c }{ m }dt}=e ^{\frac{ ct }{ m }}\]

OpenStudy (katielong):

\[\frac{ d }{ dt }(e ^{ct/m}v)=\int\limits_{}^{}g e ^{ct/m}dt\]

OpenStudy (bobo-i-bo):

Correction: \[\frac{d}{dt}(e^{ct/m}v)=ge^{\frac{ct}m}\] Which then implies that: \[e^{\frac{ct}m}v=\int \frac{d}{dt} (e^{\frac{ct}m}v) dt = \int ge^{\frac{ct}m}dt \]

OpenStudy (bobo-i-bo):

Can you solve from there?

OpenStudy (katielong):

@Bobo-i-bo I'm not too sure, I can never get the correct answer

OpenStudy (bobo-i-bo):

Can i see your attempt? Then i will be able to see where you went wrong

OpenStudy (katielong):

\[v e ^{ct/m}=\frac{ mg }{ c } e ^{ct/m}\]

OpenStudy (bobo-i-bo):

you need to add an integration constant ;)

OpenStudy (katielong):

... +C

OpenStudy (katielong):

haha, so what next?

OpenStudy (katielong):

do the exponentials cancel?

OpenStudy (bobo-i-bo):

You need to determine the "+C". But from the information you have given me... you can't determine C... so you need to give me more info!

OpenStudy (bobo-i-bo):

I'm guessing the extra info needed is that v=0 when t=0? ;P

OpenStudy (katielong):

okay, the question was: An object with mass m is dropped from rest and we assumed that air resistance is proportional to the speed of the object. If s(t) is the distance dropped after t seconds, then the speed is v=s'(t) and the acceleration a=v'(t). If g is the acceleration due to gravity, then the downward force on the object is mg-cv, where c is a positive constant and Newton II gives .... > that equation < ....

OpenStudy (katielong):

so t=0 when v=0?

OpenStudy (bobo-i-bo):

"An object with mass m is dropped from rest" :P

OpenStudy (katielong):

ahhhh so the integration constant is -mg/c

OpenStudy (bobo-i-bo):

:)

OpenStudy (katielong):

ah sneaky, thank you very much!